Find the slope of a line passing through the points: ( 2 3 , 1 5 ) and

Find the slope of a line parallel to the line in question 12.

View keyboard shortcuts12pt

Find the slope of a line perpendicular to the line from #12

View keyboard shortcuts

Paragraph
Answer #1

Slope Of line is given by : (y2-y1) /(x2-x1)

Therfore slow of line is =(5-3)/(1-2) = 2/(-1) = - 2

Slope = - 1

Since parallel lines have equal slope

Therfore slope of parallel line =-1

Since slope of perpendicular lines are negative reciprocal of each other

Therefore slope of perpendicular line = 1